# SOPHY: Learning to Generate Simulation-Ready Objects with PHYsical Materials

### v2

@@ 4.1 Shape, Texture and Material Autoencoder
As shown in Figure 4, the decoded object $\hat{\boldsymbol{X}}$ exhibits a rather flat and blurry visual appearance due to insufficient texture resolution inherited by the vertex-based color representation. Thus, we refine its texture based on renderings of the input shape (in the auto-encoding setting) or conditioned signals (in the generative setting). Specifically, we first generate a UV map, unwrapping the generated object, and then bake its per-vertex color information into a 2D texture based on the UV layout. Subsequently, we adopt the texture enhancement method of Paint3D [83] to refine the 2D texture for improved visual quality. This process serves as an optional post-processing step for our autoencoder. In general, it takes around 1 minute to refine the texture of an object.
@@ 4.2 Diffusion
Generating a simulation-ready object involves executing the reverse process of a diffusion that progressively transforms Gaussian noise in the latent space of objects into target latent codes: $\boldsymbol{Z}=\{\boldsymbol{z}_{i}\}_{i=1}^{M}$. Sampling is performed by solving the stochastic differential equations from the EDM diffusion pipeline [26]. Once a latent code is sampled, we pass it through our trained decoder to extract occupancy values on a dense $\mathbb{R}^{3}$ grid. These values are then converted into a mesh using marching cubes, then color and material properties are decoded at densely sampled mesh points.
@@
where $D$ is the denoising network, $\hat{\boldsymbol{Z}}$ are training latent codes, $\boldsymbol{N}$ are added Gaussian noises, $\sigma_{t}$ denotes the noise level, and $\boldsymbol{c}$ is a signal for conditioning. We experimented with two input conditions to our denoiser: (a) an RGB image of a target object, where $\boldsymbol{c}$ represents here the extracted features from a pre-trained image encoder (“DINOv2-ViT-B/14” [51]), (b) a text prompt, where $\boldsymbol{c}$ represents features from a pre-trained text encoder (“CLIP-ViT-L/14” [56]). The denoising network consists of alternating self-attention layers for processing the latent representations and cross-attention layers for incorporating the signals $\boldsymbol{c}$. Details on the denoiser architecture and conditional signals are provided in Section B.1 and Section B.2, respectively.
@@ 5 Experiments
MAE-$\log(E)$downarrow 1.18 0.50 0.47 0.45
MAE-$\nu$(×$\times 10^{-2}$) $\downarrow$ 4.10 3.06 2.98 3.06
MAE-$\log(\sigma)$downarrow 1.07 0.41 0.32 0.29
MAE-$\phi$ (×$\times 10^{-2}$) $\downarrow$ 5.45 1.89 1.22 1.28
MAE-$\rho$downarrow 0.16 0.14 0.13 0.09
Sim-CD (×$\times 10^{-3}$) $\downarrow$ 53.84 17.47 10.05 8.72
MAE-$c$(×$\times 10^{-2}$) $\downarrow$ 8.75 8.47 8.35 8.13
@@
Metric DG4D B. Dec. B. Perc. SOPHY
VideoPhy $\uparrow$ -0.36 -0.22 0.18 0.40
CLIP $\uparrow$ 0.71 0.73 0.74 0.77
Time (min) $\downarrow$ 10 5 7 5
@@
Metric STAG4D B. Dec. B. Perc. SOPHY
VideoPhy $\uparrow$ -1.20 -0.16 0.59 0.76
CLIP $\uparrow$ 0.13 0.15 0.18 0.18
Time (min) $\downarrow$ 66 5 7 5
@@ 5.1 Auto-encoding Evaluation
In terms of auto-encoding evaluation, our goal is to check how well we are able to recover a test 3D shape, along with its color and material properties, given its input representation. This evaluation is common in 3D latent-based generative models [85], since it is imperative for the autoencoder to capture latent spaces that can generalize to novel inputs.
@@
For material and color prediction, we report the classification accuracy of the material behavior type (M.B. Acc.) and the mean absolute error (MAE) of all our material parameters. In addition, we report Chamfer distance (Sim-CD) between densely sampled points of the predicted shapes and ground-truth ones under the deformed states computed by our MPM simulator along the whole simulation trajectory for the dropping test scenario (Section 3.2). Finally, we report purely geometric measures for the rest state of the reconstructed test shape compared to the ground truth. We use the standard metrics of Chamfer Distance (CD), volumetric Intersection-over-Union (IoU), and F-score, as used in prior work [85]. We average all values for each object in our test set.
@@ Results.
Section 5 presents quantitative comparisons in the auto-encoding setting. Compared to the baseline model, jointly modeling shape, color, and material in a shared embedding space significantly improves material metrics, with additional, more modest gains in color and geometry evaluation metrics. Notably, the material behavior type classification is improved by more than $20\%$, while Sim-CD is reduced by a factor of 5×5$\times$5 ×. This performance gain supports our hypothesis that geometry, color, and material attributes are strongly correlated and should be jointly modeled in the generative process. Compared to not using cross-attention in the decoder (“w/o CA”) model, we see that our full model has 2×2$\times$2 × lower Chamfer distance during simulation, and still slightly better performance in all other metrics. Finally, we observe that the “Fused” variant, has a tiny edge over our model in Poisson’s ratio and friction angle prediction. Yet, it performs significantly worse in terms of Sim-CD, and worse in all other measures. We suspect this behavior is due to an uneven network capacity allocation across occupancy, color, and material within a single fused latent code.
@@ 5.2 Image-to-4D
We first compare with a popular image-to-4D method, DreamGaussian4D (DG4D) [57], which generates deforming 3DGS over time conditioned on an input image, yet does not rely on any explicit physics-based representation. We also compare with our “baseline” discussed in the previous section. In addition, we compare with another baseline inspired by recent work [41, 9], which adopts a two-stage framework for simulation-ready object generation (see Section 2), and design a training-free baseline, which we call “B. Perc”. Specifically, this alternative baseline uses the perceptual models [45, 50] for material property estimation given an off-the-shelf 3D generation model [79]. After obtaining the material properties, we use the same simulator to create dynamics as SOPHY does. We provide more details about this alternative baseline in Section B.4.
@@
Given rendered images of the 4D scenes generated by any of the above compared methods, we leverage VideoPhy [3], a state-of-the-art VLM trained based on human annotations, to evaluate whether the generated videos align with real-world physics. The method produces a per-scene score that is uncalibrated for different scenarios. Thus, following [67], we apply $z$-score normalization to calibrate the scores across all methods for each scene and report the average values. A positive $z$-score indicates a method performs better than average, while negative means worse. Additionally, we use the CLIP [56] score to evaluate the consistency between the generated results and the input conditions. We also report the average time to create a video for an object starting from an image.
@@ Results.
In Section 5, SOPHY outperforms the other methods, demonstrating the highest z-score for alignment with real-world physics based on VideoPhy. For CLIP scores, our method is better aligned with the conditioned signals. We also show qualitative results for a few samples for test scenes in Figure 5. We observe that DG4D struggles to generate physically plausible animations, e.g., the headband size fluctuates over time, as it overlooks physical constraints. Both baselines, “B. Dec.” and “B. Perc.”, generate geometries and physical materials independently. Thus, they often create physically implausible objects (e.g., strange chairs in Figure 5(b)) or predict inappropriate materials for shapes (e.g., too stiff materials for a bag that looked like it was made of fabric in Figure 5(c)). In contrast, SOPHY produces more realistic dynamics, aligning closely with the given images.
@@ 5.3 Text-to-4D
We also evaluate SOPHY for text-to-4D generation. The evaluation is the same as before with the only difference being the input condition (text instead of an image). We test on prompts, describing the object category, part tags, fine-grained material categories for our test objects, following the prompts of Section 3.1. We also include the test scenario in the prompt.
@@
We compare our method with STAG4D [84], which generates deforming 3DGS driven by text prompts. We also consider the two baselines described in the previous section here. The evaluation metrics are the same as those used in the image-to-4D setting.
@@ Results.
In Section 5, it is observed that SOPHY achieves the highest performance in VideoPhy and CLIP scores, demonstrating better alignment with real-world physics and given text prompts. We also present a visual comparison for the generated dynamics in Figure 6. We observe that STAG4D fails to produce noticeable deformations, even for soft objects like teddy bears. Since the two baselines exclude material properties from the generation process, they fail to produce appropriate physical attributes aligned with the text prompts. For instance, in Figure 6(a), “B. Dec.” generates a very stiff plant while “B. Perc.” creates a soft metal container. Compared to them, our method produces more geometrically coherent and physically plausible results.
@@ 6 Conclusion
We presented a new generative model of 3D objects incorporating geometry, color, and physical material properties. Our experiments demonstrated significant benefits of the approach, including generating physically plausible 4D videos from images and text. Interesting avenues for future work include extending our generative model for synthesizing whole scenes and generating other phenomena e.g., fluids [36, 48, 69] or gases [17]. More accurate simulators, e.g., based on finite elements [62, 4, 68], could be used instead since our model generates the full objects’ volume. Our annotated dataset is only a first step towards more physics-aware 3D datasets – enlarging it with more objects and material properties would be a fruitful direction.
@@ A.1 Conditioning on images from other datasets
Here, we investigate how well SOPHY generalizes to conditioned images sourced from different datasets. We specifically select the Objaverse [16], a widely used dataset in previous 3D generation research, to obtain 2D renderings of several 3D objects that serve as input conditional signals for our method. After completing the physics-aware 3D generation, we simulate the generated assets using the pipeline described in Section 3.2 of the manuscript, finally obtaining dynamic videos. We present the results in Figure 7. Even though our method was not specifically trained on this dataset, it can still generate 3D assets with plausible material properties and physical simulations.
@@ A.2 Conditioning on text prompt variations
In this subsection, we study whether SOPHY can generate different 3D objects aligned well with the prompts based on edited text conditions. Specifically, we qualitatively demonstrate changes in the generated results by changing object parts or categories in the input prompt. The results are shown in Figure 8. Starting with the text shown in Figure 8(a), we add a description of “a polyester pillow” in (b). The resulting generations accurately reflect this change in the descriptions. Furthermore, we alter the object category from “sofa” in (a) to “chair” in (c). As shown in Figure 8(c), SOPHY can generate similar shapes for the unedited parts in the prompt, such as “cushion” and “armrest”, while the rest of the shape changes to reflect the edited category.
@@ Appendix B Implementation Details
Figure 9: Denoiser architecture. We adopt self attention (SA) and cross attention (CA) blocks to learn denoised representations starting from latents (see Equation 3). The cross attention incorporates conditions $\boldsymbol{c}$ (images/prompts).
@@ B.1 Denoiser Architecture
In Figure 9, we present the architecture of the denoising network used in our method. We alternate self-attention (SA) and cross-attention (CA) layers. The cross attention is used to inject the condition signal $\boldsymbol{c}$ (image or text prompt) into the latent representation.
@@ B.4 Perception-based Baseline
In Section 5.2, we introduced the perception-based baseline (i.e., “B. Perc.”). The baseline uses perceptual models to estimate material properties based on an off-the-shelf 3D generation model. Here, we present more details of how this baseline is implemented. Specifically, we adopt TRELLIS [79], a state-of-the-art 3D generation model, to generate textured 3D shapes given image or text conditions. To obtain the material properties of each generated shape, we then leverage an open-vocabulary 3D part segmentation model, Find3D [45], to get the part labels for the sampled surface points. Note that the query part names we provide to Find3D are derived from the set of part labels in our dataset, which comes from 3DCoMPaT200 [1]. Finally, we leverage ChatGPT-4o [50] by providing it with two renderings of the generated 3D object and asking it to estimate the material properties for each part of the object retrieved by Find3D. The prompts we employed follow the one used in our dataset collection pipeline, as described in Section D.2. After we acquire the material information for each part of the 3D object, we propagate this information to the points sampled within the object’s volume using the same strategy outlined in Section B.3. Finally, we simulate the generated object with the estimated material properties.
@@ B.5 Training Details
For object autoencoding, we use a surface point cloud of $2,048$ points as input to the autoencoder. At each iteration, we sample $1,024$ query points from the bounding sphere of the object and another $1,024$ points from the surface region for attribute prediction. Following [86], we ensure an equal distribution of query points between occupied and non-occupied regions. We set the weight coefficients described in Equation 6 as $\lambda_{c}=\lambda_{\nu}=\lambda_{\phi}=\lambda_{\rho}=\lambda_{M}=0.1$, $\lambda_{E}=\lambda_{\sigma}=0.01$, and $\lambda_{r}=0.001$. The autoencoder is trained with a batch size of $256$ for $500$ epochs, using a learning rate that linearly increases to $10^{-4}$ over the first $25$ epochs before gradually decaying to $10^{-5}$ following a cosine schedule.

### v3

@@ 4.1 Shape, Texture and Material Autoencoder
As shown in Figure˜4, the decoded object $\hat{\boldsymbol{X}}$ exhibits a rather flat and blurry visual appearance due to insufficient texture resolution inherited by the vertex-based color representation. Thus, we refine its texture based on renderings of the input shape (in the auto-encoding setting) or conditioned signals (in the generative setting). Specifically, we first generate a UV map, unwrapping the generated object, and then bake its per-vertex color information into a 2D texture based on the UV layout. Subsequently, we adopt the texture enhancement method of Paint3D zeng2024paint3d to refine the 2D texture for improved visual quality. This process serves as an optional post-processing step for our autoencoder. In general, it takes around 1 minute to refine the texture of an object.
@@ 4.2 Diffusion
Generating a simulation-ready object involves executing the reverse process of a diffusion that progressively transforms Gaussian noise in the latent space of objects into target latent codes: $\boldsymbol{Z}=\{\boldsymbol{z}_{i}\}_{i=1}^{M}$. Sampling is performed by solving the stochastic differential equations from the EDM diffusion pipeline karras2022elucidating . Once a latent code is sampled, we pass it through our trained decoder to extract occupancy values on a dense $\mathbb{R}^{3}$ grid. These values are then converted into a mesh using marching cubes, then color and material properties are decoded at densely sampled mesh points.
@@
where $D$ is the denoising network, $\hat{\boldsymbol{Z}}$ are training latent codes, $\boldsymbol{N}$ are added Gaussian noises, $\sigma_{t}$ denotes the noise level, and $\boldsymbol{c}$ is a signal for conditioning. We experimented with two input conditions to our denoiser: (a) an RGB image of a target object, where $\boldsymbol{c}$ represents here the extracted features from a pre-trained image encoder (“DINOv2-ViT-B/14” oquab2024dinov ), (b) a text prompt, where $\boldsymbol{c}$ represents features from a pre-trained text encoder (“CLIP-ViT-L/14” radford2021learning ). The denoising network consists of alternating self-attention layers for processing the latent representations and cross-attention layers for incorporating the signals $\boldsymbol{c}$. Details on the denoiser architecture and conditional signals are provided in Section˜B.1 and Section˜B.2, respectively.
@@ 5 Experiments
MAE-$\log(E)$ $\downarrow$ 1.18 0.50 0.47 0.45
MAE-$\nu$($\times 10^{-2}$) $\downarrow$ 4.10 3.06 2.98 3.06
MAE-$\log(\sigma)$ $\downarrow$ 1.07 0.41 0.32 0.29
MAE-$\phi$ ($\times 10^{-2}$) $\downarrow$ 5.45 1.89 1.22 1.28
MAE-$\rho$ $\downarrow$ 0.16 0.14 0.13 0.09
Sim-CD ($\times 10^{-3}$) $\downarrow$ 53.84 17.47 10.05 8.72
MAE-cc($\times 10^{-2}$) $\downarrow$ 8.75 8.47 8.35 8.13
@@
Metric DG4D Free4D B. Dec. B. Perc. SOPHY
VideoPhy $\uparrow$ -0.28 -0.21 -0.15 0.21 0.43
CLIP $\uparrow$ 0.71 0.74 0.73 0.74 0.77
Time (min) $\downarrow$ 10 73 5 7 5
@@
Metric STAG4D Free4D B. Dec. B. Perc. SOPHY
VideoPhy $\uparrow$ -1.23 0.07 -0.18 0.58 0.76
CLIP $\uparrow$ 0.13 0.17 0.15 0.18 0.18
Time (min) $\downarrow$ 66 73 5 7 5
@@ 5.1 Auto-encoding Evaluation
In terms of auto-encoding evaluation, our goal is to check how well we are able to recover a test 3D shape, along with its color and material properties, given its input representation. This evaluation is common in 3D latent-based generative models zhang20233dshape2vecset , since it is imperative for the autoencoder to capture latent spaces that can generalize to novel inputs.
@@
For material and color prediction, we report the classification accuracy of the material behavior type (M.B. Acc.) and the mean absolute error (MAE) of all our material parameters. In addition, we report Chamfer distance (Sim-CD) between densely sampled points of the predicted shapes and ground-truth ones under the deformed states computed by our MPM simulator along the whole simulation trajectory for the dropping test scenario (Section˜3.2). Finally, we report purely geometric measures for the rest state of the reconstructed test shape compared to the ground truth. We use the standard metrics of Chamfer Distance (CD), volumetric Intersection-over-Union (IoU), and F-score, as used in prior work zhang20233dshape2vecset . We average all values for each object in our test set.
@@ Results.
Section˜5 presents quantitative comparisons in the auto-encoding setting. Compared to the baseline model, jointly modeling shape, color, and material in a shared embedding space significantly improves material metrics, with additional, more modest gains in color and geometry evaluation metrics. Notably, the material behavior type classification is improved by more than $20\%$, while Sim-CD is reduced by a factor of $5\times$. This performance gain supports our hypothesis that geometry, color, and material attributes are strongly correlated and should be jointly modeled in the generative process. Compared to not using cross-attention in the decoder (“w/o CA”) model, we see that our full model has $2\times$ lower Chamfer distance during simulation, and still slightly better performance in all other metrics. Finally, we observe that the “Fused” variant, has a tiny edge over our model in Poisson’s ratio and friction angle prediction. Yet, it performs significantly worse in terms of Sim-CD, and worse in all other measures. We suspect this behavior is due to an uneven network capacity allocation across occupancy, color, and material within a single fused latent code.
@@ 5.2 Image-to-4D
We first compare SOPHY with conventional image-to-4D methods, which generate deforming 3DGS over time without any explicit physics-based representation. We examine DreamGaussian4D (DG4D) ren2023dreamgaussian4d , a pioneering work in this field, along with Free4D liu2025free4d , which is one of the state-of-the-art methods that distills pre-trained foundation models for consistent 4D generation. We also compare with our “baseline” discussed in the previous section. In addition, we consider another physics-aware baseline inspired by recent work lin2024phys4dgen ; chen2025physgen3d , which adopts a two-stage framework for simulation-ready object generation (see Section˜2), and design a training-free baseline, which we call “B. Perc”. Specifically, this alternative baseline uses the perceptual models ma2024find ; hurst2024gpt4o for material property estimation given an off-the-shelf 3D generation model xiang2024structured . After obtaining the material properties, we use the same simulator to create dynamics as SOPHY does. We provide more details about this alternative baseline in Section˜B.4.
@@
Given rendered images of the 4D scenes generated by any of the above compared methods, we leverage VideoPhy bansal2024videophy , a state-of-the-art VLM trained based on human annotations, to evaluate whether the generated videos align with real-world physics. The method produces a per-scene score that is uncalibrated for different scenarios. Thus, following tan2024physmotion , we apply $z$-score normalization to calibrate the scores across all methods for each scene and report the average values. A positive $z$-score indicates a method performs better than average, while negative means worse. Additionally, we use the CLIP radford2021learning score to evaluate the consistency between the generated results and the input conditions. We also report the average time to create a video for an object starting from an image.
@@ Results.
In Section˜5, SOPHY outperforms the other methods, demonstrating the highest z-score for alignment with real-world physics based on VideoPhy. For CLIP scores, our method is better aligned with the conditioned signals. We also show qualitative results for a few samples for test scenes in Figure˜5. We observe that DG4D and Free4D struggle to generate physically plausible animations, e.g., the headband size fluctuates over time, as they overlook physical constraints. Both baselines, “B. Dec.” and “B. Perc.”, generate geometries and physical materials independently. Thus, they often create physically implausible objects (e.g., strange chairs in Figure˜5(b)) or predict inappropriate materials for shapes (e.g., too stiff materials for a bag that looked like it was made of fabric in Figure˜5(c)). In contrast, SOPHY produces more realistic dynamics, aligning closely with the given images.
@@ 5.3 Text-to-4D
In this subsection, we evaluate SOPHY for text-to-4D generation. The evaluation is the same as before, with the only difference being the input condition (text instead of an image). We test on prompts that describe the object category, part tags, and fine-grained material categories for our test objects, following the prompts of Section 3.1. We also include the test scenario in the prompt.
@@
We compare our method with STAG4D zeng2024stag4d , which generates deforming 3DGS driven by text prompts. We also consider Free4D liu2025free4d and the two baselines in the previous section here. The evaluation metrics are the same as in the image-to-4D setting.
@@ Results.
In Section˜5, it is observed that SOPHY achieves the highest performance in VideoPhy and CLIP scores, demonstrating better alignment with real-world physics and given text prompts. We also present a visual comparison for the generated dynamics in Figure˜6. We observe that STAG4D fails to produce noticeable deformations, even for soft objects like teddy bears. Although Free4D can occasionally generate dynamic content as shown in Figure˜6(b), the content mismatches the conditional signal. For the two baselines, since they exclude material properties from the generation process, they fail to produce appropriate physical attributes aligned with the text prompts. For instance, in Figure˜6(a), “B. Dec.” generates a very stiff plant while “B. Perc.” creates a soft metal container. Compared to them, our method produces more geometrically coherent and physically plausible results.
@@ 6 Conclusion
We presented a new generative model of 3D objects incorporating geometry, color, and physical material properties. Our experiments demonstrated significant benefits of the approach, including generating physically plausible 4D videos from images and text. Interesting avenues for future work include extending our generative model for synthesizing whole scenes and generating other phenomena e.g., fluids li2024neural ; muller2003particle ; ummenhofer2019lagrangian or gases feldman2005animating . More accurate simulators, e.g., based on finite elements sifakis2012fem ; vega ; trusty2022mixed , could be used instead since our model generates the full objects’ volume. Our annotated dataset is only a first step towards more physics-aware 3D datasets – enlarging it with more objects and material properties would be a fruitful direction.
@@ A.1 Conditioning on images from other datasets
Here, we investigate how well SOPHY generalizes to conditioned images sourced from different datasets. We examine both synthetic and real-world scanned objects. Specifically, we choose the Objaverse [16], a widely used synthetic dataset in previous 3D generation research, along with the Google Scanned Objects [17], which is a high-quality dataset featuring scanned everyday items. We obtain 2D renderings of several 3D objects from them to serve as input conditional signals for our method. After completing the physics-aware 3D generation, we simulate the generated assets using the pipeline described in Section˜3.2 of the manuscript, finally obtaining dynamic videos. We present the results in Figure˜7 and Figure˜8. Even though our method was not specifically trained on these datasets, it can still generate 3D assets with plausible material properties and physical simulations.
@@
Figure 8: Generated results by conditioning our method on images from Google Scanned Objects [17] shown on the left.
@@ A.2 Conditioning on text prompt variations
In this subsection, we study whether SOPHY can generate different 3D objects aligned well with the prompts based on edited text conditions. Specifically, we qualitatively demonstrate changes in the generated results by changing object parts or categories in the input prompt. The results are shown in Figure˜9. Starting with the text shown in Figure˜9(a), we add a description of “a polyester pillow” in (b). The resulting generations accurately reflect this change in the descriptions. Furthermore, we alter the object category from “sofa” in (a) to “chair” in (c). As shown in Figure˜9(c), SOPHY can generate similar shapes for the unedited parts in the prompt, such as “cushion” and “armrest”, while the rest of the shape changes to reflect the edited category.
@@ Appendix B Implementation Details
Figure 10: Denoiser architecture. We adopt self attention (SA) and cross attention (CA) blocks to learn denoised representations starting from latents (see Equation˜3). The cross attention incorporates conditions $\boldsymbol{c}$ (images/prompts).
@@ B.1 Denoiser Architecture
In Figure˜10, we present the architecture of the denoising network used in our method. We alternate self-attention (SA) and cross-attention (CA) layers. The cross attention is used to inject the condition signal $\boldsymbol{c}$ (image or text prompt) into the latent representation.
@@ B.4 Perception-based Baseline
In Section˜5.2, we introduced the perception-based baseline (i.e., “B. Perc.”). The baseline uses perceptual models to estimate material properties based on an off-the-shelf 3D generation model. Here, we present more details of how this baseline is implemented. Specifically, we adopt TRELLIS [85], a state-of-the-art 3D generation model, to generate textured 3D shapes given image or text conditions. To obtain the material properties of each generated shape, we then leverage an open-vocabulary 3D part segmentation model, Find3D [50], to get the part labels for the sampled surface points. Note that the query part names we provide to Find3D are derived from the set of part labels in our dataset, which comes from 3DCoMPaT200 [1]. Finally, we leverage ChatGPT-4o [55] by providing it with two renderings of the generated 3D object and asking it to estimate the material properties for each part of the object retrieved by Find3D. The prompts we employed follow the one used in our dataset collection pipeline, as described in Section˜D.2. After we acquire the material information for each part of the 3D object, we propagate this information to the points sampled within the object’s volume using the same strategy outlined in Section˜B.3. Finally, we simulate the generated object with the estimated material properties.
@@ B.5 Training Details
Figure 11: Statistics of fine-grained material types in our dataset.
For object autoencoding, we use a surface point cloud of $2,048$ points as input to the autoencoder. At each iteration, we sample $1,024$ query points from the bounding sphere of the object and another $1,024$ points from the surface region for attribute prediction. Following [92], we ensure an equal distribution of query points between occupied and non-occupied regions. We set the weight coefficients described in Equation˜6 as $\lambda_{c}=\lambda_{\nu}=\lambda_{\phi}=\lambda_{\rho}=\lambda_{M}=0.1$, $\lambda_{E}=\lambda_{\sigma}=0.01$, and $\lambda_{r}=0.001$. The autoencoder is trained with a batch size of $256$ for $500$ epochs, using a learning rate that linearly increases to $10^{-4}$ over the first $25$ epochs before gradually decaying to $10^{-5}$ following a cosine schedule.